OpenStudy (anonymous):

Multiply and simplify. Assume that all variables represent nonnegative numbers. (square root 6+3)( square root 6-1)

OpenStudy (tkhunny):

What an odd thing to say. There are no variables. What's preventing you from multiplying these binomials? Can you do these: (x+2)(x-4)?

OpenStudy (anonymous):

(square root 9)(square root 5)

OpenStudy (tkhunny):

What is that? That has nothing to do with either the original problem statement or my inquisitive example. If you show your intermediate steps - your work - we can see where you wander off.

OpenStudy (anonymous):

im lost

OpenStudy (tkhunny):

Okay, answer my question. CAn you multiply these binomials? (x+2)(x-4)

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Please demonstrate. Do NOT just write the final result.

OpenStudy (anonymous):

x*x= x, 2*4=8

OpenStudy (tkhunny):

Only two pieces? That is not correct. Okay, try this. (10 + 3)(10 + 4)

OpenStudy (anonymous):

know I can multiply them right

OpenStudy (tkhunny):

Notice how in multiplying 13 * 14 there are FOUR pieces to consider. (10 + 3)(10 + 4) = 10*10 + 10*4 + 3*10 + 3*4 Do you see all four of those pieces?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Okay, now to my first question. What are the FOUR pieces of this multiplication? (x+2)(x-4)

OpenStudy (anonymous):

(10*10)(10*4)(3*10)(3*4)

OpenStudy (tkhunny):

Well, that's a good enough LIST, but it is not a good enough result. They should be connected by addition. Have you met the "Distributive Property"?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

That is ALL we are doing, here. (10 + 3)(10 + 4) = 10(10 + 4) + 3(10 + 4) = 10*10 + 10*4 + 3*10 + 3*4 It's all the Distributive Property - Nothing else.

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

Okay, let's move on to the other example. Multiply these. Use the Distributive Property to find the FOUR pieces. (x+2)(x-4)

OpenStudy (anonymous):

(x+2)(x-4)=x(x+2)+2(x-4)=x*x+x*4+2*x+2*4

OpenStudy (tkhunny):

Not bad, but it is a little confused. Look at this and spot the difference... (x+2)(x-4) = x(x-4) + 2(x-4) <== See how it is (x-4) both times? This is important. (x+2)(x-4) = x(x-4) + 2(x-4) = x*x - x*4 + 2*x - 2*4 <== This one is additionally different because I noticed the negative signs. All yours turned into "+"! That's no good.